OpenStudy (anonymous):

Write an equation in point-slope form, slope-intercept form, and standard from for these conditions. slope 3/2, y-intercept 4

OpenStudy (luigi0210):

Point slope: \[y-y_{1}=m(x-x_{1})\] Standard form: \[Ax+By=C\] Slope Intercept form: \[y=mx+b\]

OpenStudy (luigi0210):

m=slope x1=x coordinate y1=y coordinate

OpenStudy (anonymous):

i know the formula's but how do you start when your given a y-intercept rather then an ordered pair??

OpenStudy (luigi0210):

y-intercept is when x=0

OpenStudy (luigi0210):

So (0, 4)

OpenStudy (anonymous):

oh okay so then i do y-4=3/2(x-0)

OpenStudy (anonymous):

is this correct for point slope y-4=3/2(x-0)

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

then for slope-intercept Form y=3/2x+4

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

then for standard form its 3/2x+y=4 A=3/2 B=1 C=4

OpenStudy (luigi0210):

You can't have a fraction in front of the A, so multiply it by it's reciprocal

OpenStudy (anonymous):

so A should = 1? 3/2 x 2/3= 6/6=1?

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

thanks man :D

OpenStudy (luigi0210):

Anytime :)